If the nominal rate of interest is 13.82% and the real rate of interest is 8.03%​, what is the expected rate of​ inflation?
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.0803)=(1+0.1382)/(1+inflation rate)

(1+inflation rate)=(1+0.1382)/(1+0.0803)

inflation rate=[(1.1382)/(1.0803)]-1

=5.36%(Approx)